User Tools

Site Tools


lbaops:datamanagement:oct10disktransfer

Calendar of Black-listed times

Note maximum transfer speeds:
Hobart MUST be below 100Mbps.
Parkes needs plenty of bandwidth spare for pulsar data. See Calendar below.
ATCA+Mp+Parkes must be <512Mbps.

See also individual telescope schedules:
Hobart November 2010 schedule
Parkes Oct10-Mar11 schedule
ATCA 25th Oct - 7th Nov
Mopra 25th Oct - 7th November

(:includeurl Pmcal.Pmcal height=600px :)

Data Transfer to Pbstore

Can check these against the full transfers summary page.

Experiment Disk Location Status Received
Hobart (hovsi)
vc124 ATNF V009A Done 128 - ok
v186b ATNF V009A hovsi:/xraid/Br_1 Done Ok
v255l ATNF V009A Done Ok
v316d ATNF V016B Done Ok
v255m ATNF V016B Done Ok
v190o CURT V003B hovsi:/xraid/Al_1 Done Ok
v329b CURT V003B Done Ok
v252t CURT V001B, UTAS V002A Done Ok
v252u UTAS V001A Done Ok
Parkes
v186b ATNF V010B pkvsi1:xraid0 Done Ok
v255l ATNF V001B pkvsi1:/xraid1 + 83 files on /xraid0 Done Ok
v316d ATNF V001B, ATNF V010B pkvsi1:/xraid1, pkvsi1:/xraid0 Done Ok
v255m ATNF V014A pkvsi2:xraid0 Done Ok
v190o ATNF V014A pkvsi2:xraid0 Done Ok
v329b ATNF V013B pkvsi2:/xraid1 Done Ok
v252t ATNF V014B pkvsi1:/xraid0 Done Ok
v252u CURT V002B pkvsi1:/xraid1 Done Ok
vx016c pkvsi1:/xraid0 Done Ok
ATCA
v186b CURT V002A cavsi2:xraid0 Done Ok
v255l ATNF V002A cavsi2:xraid1 Done ok
v316d ATNF V002A, CURT V002A cavsi2:xraid1,xraid0 Done Ok
v255m ATNF V007B Done Ok
v190o ATNF V007B Done Ok
v329b ATNF V008A Done Ok
v252t CAVE cave-store:/data2 Done Ok
v271g ATNF V009B, ATNF V005A cavsi1:/xraid0, /xraid1 Done OK
v252u CAVE cave:/data2 Done Ok
vx016c cavsi2:/xraid0 Done Ok
Mopra
v186b CAVE cave:/data3 Done OK
v255l CAVE cave:/data1 Done Ok
v316d CAVE cave:/data3 Done Ok
v255m CAVE cave-store:/data1 Done Ok
v190o CAVE cave:/data1 Done Ok
v329b CAVE cave:/data2 Done Ok
v252t ATNF V004B mpvsi2:xraid0 Done Ok
v271g ATNF V002B, ATNF V015A mpvsi1:/xraid1, /xraid0 Done Ok
v252u UTAS V001B, ATNF V007B mpvsi2:/xraid1, cavsi2:/xraid1 Done Ok
vx016c CAVE cave:/data1 Done 4 - ok
Tid
v190o ATNF V011/A1-7 → ATNF-E001 2070 - ok
v252t ATNF V011/A1-7 → ATNF-E001/E002 2712 - ok
v329b ATNF V011/A1-7 → ATNF-E001 318 - Ok
v252u ATNF V011/A1-7 → ATNF-E003 966 - Ok
vt16h ATNF V011/B1-7 → ATNF-E003/E001 1179 - Ok
Hart - CHRIS TO HANDLE
v316d HART-0051/4000/1024
v255l HART-0051/4000/1024
v255m HART-0051/4000/1024
v252t HART-0051/4000/1024 1 file to be recopied
v252u HART-0052/4000/1024 Ok
lbaops/datamanagement/oct10disktransfer.txt · Last modified: 2015/12/18 16:38 (external edit)